OpenStudy (anonymous):

Container A holds 752 mL of ideal gas at 2.10 atm. Container B holds 184 mL of ideal gas at 4.80 atm. If the gases are allowed to mix together, what is the resulting pressure?

OpenStudy (cuanchi):

There are at least two different ways to calculate the final pressure, One is calculate how many moles of A and the moles of B that you have with the equation PV=nRT n=PV/RT R=0.082 atm L/ mol K T= any value different than 0 Add the two values (moles of A+ moles of B= total moles of gas) Add the two volumes (mL of A+ mL of B= total mL of gas = total mLx10^(-3) L gas convert to L to be able to use R Finally calculate the final pressure with the same formula PV=nRT => P= nRT/V n will be now the total moles of gas and V the total volume of gas The other way is to do the formula calculation and eliminate R and T \[P _{final}= \frac{ P _{A} \times V _{A} + P _{B} \times V _{B} }{ V _{A}+ V _{B} }\] Explanation \[n _{A}=\frac{ P _{A} \times V _{A}}{ R \times T } \] \[n _{B}=\frac{ P _{B} \times V _{B}}{ R \times T } \] \[n _{total}= n _{A}+ n _{B}\] \[n _{total}= \frac{ P _{A} \times V _{A}}{ R \times T } + \frac{ P _{B} \times V _{B}}{ R \times T } \] \[n _{total}=\frac{ 1 }{R \times T } \left( P _{A} \times V _{A} +P _{B} \times V _{B} \right)\] \[P _{final}= n _{total} \times R \times T \div V _{total}\] \[V _{total}= V _{A}+V _{B}\] \[P _{final}= \frac{ 1 }{R \times T } \left( P _{A} \times V _{A} +P _{B} \times V _{B} \right) \times R \times T \div V _{total}\] cancel R x T final formula to calculate the final pressure \[P _{final}= \frac{ P _{A} \times V _{A} + P _{B} \times V _{B} }{ V _{A}+ V _{B} }\]